# Broad-Line AGN at : The Black Hole Mass Function and a Connection with Little Red Dots

### v1

@@ 3.2 Flux Calibration
While NIRSpec MSA flux calibration has improved significantly since the first data releases and pipelines (e.g. Schaerer et al., 2022; Taylor et al., 2022; Carnall et al., 2023; Trump et al., 2023), additional corrections to account for slit-losses/path-losses and overall photometric calibrations are still necessary. We use the NIRCam photometry described in §2.1 to account for these effects in our dataset. To do this, we first visibly inspect the 2D spectrum of each object for contamination. The spectra of some objects are contaminated by either positive flux from other sources in the same MSA row, or negative flux resulting from nodded background subtraction of sources in adjacent MSA rows. We visually inspect our sources and flag those that are impacted by either type of contamination. We find eight sources with significant contamination, and 42 sources that are contamination free.
@@
For contamination-free sources, we apply a simple flux correction based on NIRCam photometry. For each source, we use the NIRCam filter transmission curves and G395M spectrum to measure synthetic photometry from the G395M spectrum for both the F356W and F444W NIRCam filters. We then select the filter that overlaps the detected broad-line and scale the flux and error spectra such that its broadband spectroscopic flux density matches the flux density of the broadband photometry. While using the higher S/N PRISM spectra (when available) would be preferable for this calibration, due to the different MSA configurations between the PRISM and G395M data, the measured fluxes between the two are not directly comparable. For the remaining eight contaminated sources and three sources that fall outside the CEERS NIRCam footprint (see Figure 6), we instead apply the median corrective factor (1.28) from the 39 uncontaminated sources with NIRCam coverage to avoid any extreme calibration errors from the contamination.
After this correction, we re-fit the broad-lines using the emcee and above-mentioned broad-line model to produce final, flux-calibrated line fits and calibrated line measurements using the full calibrated line fit posterior distributions.
@@
We apply a similar procedure to calibrate the PRISM spectra (when available) of the broad-line sample. Here, for a given source, we use the transmission curves of the F150W, F200W, F277W, F365W, F410M, and F444W filters to measure broadband fluxes for each filter. We then fit a first order polynomial to the ratio of the fluxes measured from the spectra and the fluxes reported in the broadband photometry as a function of the logarithm of the filter centers. We use this best fit line as a wavelength dependant multiplicative correction curve, and apply this correction to the spectral data. We repeat this process five times to ensure reasonable convergence between the spectrum and measured photometry. We use these calibrated PRISM spectra in §5 to measure optical continuum slopes.
@@ 3.3 Outflows
We next test our sample of broad-line objects for evidence of inflows/outflows to help distinguish true BLAGN from objects that have broad emission components due to strong kinematics in the ISM. For the 19 objects at redshifts , such that the [O iii]4959,5007 doublet and H lines fall within the G395M spectra, we examine the [O iii]4959,5007 doublet for evidence of broad lines.
@@
Here, we design two models for MCMC fitting. In the first model (the outflow model), we model H and [O iii]4959,5007 each as two Gaussians (one narrow and one broad). We enforce a common FWHM for all three narrow components, and we require a common FWHM for the broad components of the [O iii]4959,5007 doublet. For each line, we require that the centers of the broad and narrow Gaussians are the same. We fix the amplitude ratio of the [O iii] doublet lines to 1:2.98 for both the narrow and broad components, and assume a common redshift and no velocity offset from the H line. We also include a simple continuum offset parameter. In the second model (the BLAGN model), we exclude the broad components of the [O iii]4959,5007 models but keep the expected broad H component (based on the broad H emission).
@@
We fit the spectra with both the outflow and BLAGN models each using 32 walkers and a 5000 step chain, of which we discard the first 3333 steps. Here, we also compute the Bayesian Information Criterion (BIC) for each model fit: , where is the number of free parameters in a fitted model, is the maximized likelihood function, and is the number of fitted data points. For 18/19 objects, the BLAGN model is preferred with BICBIC, indicating no evidence of a strong outflow. The final object (RUBIES-EGS-50052) shows BICBIC, indicating significant broad [O iii]4959,5007 emission likely caused by a strong outflow. We show the single and double component fits to this object in Figure 3. This object exhibits an [O iii]4959,5007 broad component with a FWHM of km s-1, an H broad component with a FWHM of km s-1, and an H broad component with a FWHM of km s-1. These FWHMs demonstrate the H and H lines have consistent velocity profiles indicative of a BLAGN, while the [O iii]4959,5007 double has a slightly lower velocity broad component, possibly due to kinematics independent of the BLAGN. Curiously, RUBIES-EGS-50052 is the source that was observed by both RUBIES and CEERS (as CEERS-2782). Its CEERS spectrum shows no trace of an outflow, but due to the aforementioned expected slit-losses for this source in CEERS, we expect that the absence of detected outflows may simply be due to the lower S/N of the object in the CEERS data relative to the RUBIES data. Alternatively, the outflow component may be spatially dependent, and may be be better sampled by the RUBIES shutter than the CEERS shutter. While the exact nature of these spatially dependent effects is beyond the scope of this work, the broad H and H emission from this source are clear BLAGN signatures, and we therefore keep source RUBIES-EGS-50052 in our BLAGN sample.
@@
Though we cannot perform this analysis for objects in our sample, the 19/19 objects that show clear BLAGN signatures that kinematically dominate any outflow velocity structure suggest that our full sample is likely significantly dominated by BLAGN, with low likelihood of containing a significant number of outflow dominated objects whose broad H emission is not driven by a central AGN. This result is consistent with the analysis performed in other collections of broad-line AGN detected by JWST at . Notably, Maiolino et al. (2024) also finds that outflows play a negligible role in broadening the lines of the Type-I AGN used in their analysis.
@@
Beyond outflows, Baggen et al. (2024) propose that the broad-lines seen in high-redshift galaxies with JWST may instead be generated by highly-compact galaxies with intrinsically high virial velocities that can mimic the broad-lines typically associated with BLAGN. However, as 49/50 objects in our sample show both broad and narrow components in H, we expect that these are indeed BLAGN showing narrow emission from galaxy kinematics (or the AGN narrow line regions) and broad emission from the AGN, and are not the exotically dense galaxies proposed by Baggen et al. (2024). While the final object (RUBIES-EGS-49140) is studied in Baggen et al. (2024) as a candidate for non-AGN broad-lines, we nonetheless keep this object in our BLAGN sample as the Baggen et al. (2024) model does not preclude this object from being a BLAGN, it simply offers a novel alternate explanation for its broad spectral features.
@@ 4 The BLAGN Sample
Here, we present our final sample of 50 H spectroscopically confirmed BLAGN. We show galleries of the fitted H broad-lines in Figure 4 and Figure 5, and detail the properties of these objects in Table 1. We show the positions of the BLAGN in the CEERS (EGS) and PRIMER-UDS fields, as well as the footprints of the CEERS and RUBIES NIRSpec MSA pointings in Figure 6. Three sources that lack NIRCam imaging coverage are visible in the upper left of the top panel and were targeted in the CEERS survey based on archival HST CANDELS imaging (Grogin et al., 2011; Koekemoer et al., 2011). The sample clearly exhibits a large diversity in H line shapes, [N II]6548,6583 line strengths, and redshifts. We analyze the properties of the sample below.
Source ID | R.A. | Decl. | FWHMbroad H | |||||
[deg] | [deg] | [ erg s-1 cm-2] | [km s-1] | |||||
RUBIES-UDS-44043 | 34.241817 | -5.239436 | 3.499 | 31.46 | 8.31 | 23.13 | 2710 | 7.66 |
RUBIES-EGS-50812 | 214.845487 | 52.848281 | 3.519 | 12.29 | 7.23 | 5.07 | 1747 | 6.96 |
RUBIES-UDS-21944 | 34.469218 | -5.283563 | 3.526 | 12.78 | 2.65 | 10.18 | 1340 | 6.86 |
RUBIES-UDS-46566 | 34.293543 | -5.234588 | 3.538 | 124.60 | 86.41 | 38.28 | 747 | 6.61 |
RUBIES-UDS-154183 | 34.410749 | -5.129664 | 3.546 | 59.20 | 11.99 | 47.28 | 2352 | 7.68 |
RUBIES-UDS-153207 | 34.493112 | -5.130999 | 3.597 | 316.60 | 246.25 | 72.67 | 1387 | 7.30 |
RUBIES-EGS-58237 | 214.850571 | 52.866030 | 3.651 | 59.66 | 19.99 | 39.58 | 2333 | 7.65 |
RUBIES-UDS-61627 | 34.238394 | -5.205775 | 3.654 | 7.85 | 5.09 | 2.78 | 1533 | 6.73 |
RUBIES-UDS-5496 | 34.405872 | -5.312951 | 3.655 | 12.91 | 7.82 | 5.15 | 2309 | 7.23 |
RUBIES-EGS-15825 | 215.079264 | 52.934252 | 3.666 | 71.95 | 7.25 | 64.51 | 3326 | 8.07 |
RUBIES-UDS-46885 | 34.291666 | -5.233788 | 3.730 | 25.66 | 17.98 | 7.67 | 707 | 6.26 |
RUBIES-EGS-34978 | 214.861690 | 52.818438 | 3.772 | 90.98 | 67.74 | 23.26 | 1442 | 7.13 |
RUBIES-UDS-10036 | 34.381671 | -5.303742 | 3.806 | 9.53 | 4.19 | 5.35 | 1504 | 6.87 |
RUBIES-EGS-37032 | 214.849388 | 52.811824 | 3.850 | 17.83 | 5.82 | 12.04 | 2033 | 7.31 |
CEERS-11728 | 215.084870 | 52.970738 | 3.869 | 9.80 | 7.95 | 1.89 | 1383 | 6.59 |
RUBIES-UDS-22304 | 34.399170 | -5.283007 | 3.906 | 24.15 | 20.87 | 3.32 | 912 | 6.34 |
RUBIES-UDS-16053 | 34.367104 | -5.293524 | 3.952 | 8.36 | 3.07 | 5.33 | 1372 | 6.80 |
RUBIES-UDS-11721 | 34.411039 | -5.300780 | 3.978 | 33.66 | 23.62 | 10.06 | 1894 | 7.23 |
RUBIES-UDS-8895 | 34.363041 | -5.306108 | 3.982 | 120.04 | 66.98 | 53.19 | 1487 | 7.35 |
RUBIES-UDS-30969 | 34.296356 | -5.268672 | 4.000 | 51.11 | 39.96 | 11.19 | 1023 | 6.70 |
RUBIES-UDS-44952 | 34.248349 | -5.237646 | 4.094 | 18.25 | 13.14 | 5.15 | 1412 | 6.84 |
RUBIES-UDS-31747 | 34.223757 | -5.260245 | 4.130 | 16.85 | 6.09 | 10.79 | 1924 | 7.27 |
RUBIES-EGS-28812 | 214.924149 | 52.849050 | 4.223 | 14.28 | 4.19 | 10.10 | 2115 | 7.36 |
RUBIES-UDS-35974 | 34.331644 | -5.260593 | 4.367 | 41.26 | 33.15 | 8.18 | 1134 | 6.77 |
RUBIES-UDS-63139 | 34.230848 | -5.202607 | 4.435 | 5.43 | 2.67 | 2.77 | 1072 | 6.51 |
RUBIES-UDS-48507 | 34.284578 | -5.230702 | 4.468 | 19.66 | 12.50 | 7.17 | 776 | 6.42 |
CEERS-1244 | 215.240652 | 53.036041 | 4.478 | 50.53 | 16.53 | 34.03 | 2606 | 7.82 |
CEERS-1665 | 215.178197 | 53.059349 | 4.483 | 34.27 | 27.61 | 6.70 | 777 | 6.40 |
RUBIES-EGS-29489 | 215.022071 | 52.920786 | 4.543 | 13.55 | 2.54 | 11.04 | 2120 | 7.41 |
RUBIES-UDS-19484 | 34.232426 | -5.280654 | 4.656 | 3.80 | 1.82 | 1.99 | 1319 | 6.65 |
RUBIES-EGS-45599 | 214.831103 | 52.824124 | 4.862 | 7.52 | 4.64 | 2.89 | 1463 | 6.84 |
RUBIES-EGS-6411 | 215.109185 | 52.939770 | 4.880 | 7.33 | 5.73 | 1.61 | 1568 | 6.78 |
RUBIES-EGS-42232 | 214.886792 | 52.855381 | 4.954 | 23.67 | 4.79 | 18.91 | 2410 | 7.68 |
RUBIES-EGS-46985 | 214.805654 | 52.809497 | 4.963 | 16.68 | 10.31 | 6.38 | 1661 | 7.12 |
RUBIES-EGS-17146 | 214.949482 | 52.845415 | 5.000 | 36.71 | 21.58 | 15.16 | 742 | 6.58 |
RUBIES-EGS-50052* | 214.823454 | 52.830277 | 5.240 | 37.86 | 20.21 | 17.66 | 2135 | 7.58 |
CEERS-2782* | 214.823453 | 52.830281 | 5.240 | 26.56 | 14.71 | 11.88 | 1510 | 7.19 |
RUBIES-EGS-13872 | 215.132933 | 52.970705 | 5.261 | 7.84 | 4.85 | 2.99 | 1540 | 6.93 |
RUBIES-EGS-42046 | 214.795368 | 52.788847 | 5.279 | 145.15 | 43.92 | 101.24 | 3574 | 8.40 |
RUBIES-EGS-926125 | 215.137081 | 52.988554 | 5.284 | 14.46 | 4.42 | 10.03 | 1871 | 7.35 |
RUBIES-EGS-60935 | 214.923373 | 52.925593 | 5.287 | 22.78 | 6.86 | 15.94 | 2139 | 7.57 |
RUBIES-UDS-29813 | 34.453355 | -5.270717 | 5.440 | 12.04 | 4.71 | 7.31 | 2042 | 7.38 |
CEERS-746 | 214.809142 | 52.868484 | 5.623 | 6.67 | 1.97 | 4.74 | 2055 | 7.31 |
CEERS-672 | 214.889677 | 52.832977 | 5.666 | 2.27 | 0.98 | 1.29 | 1335 | 6.66 |
RUBIES-UDS-19521 | 34.383672 | -5.287732 | 5.669 | 10.19 | 5.35 | 4.83 | 1455 | 7.01 |
RUBIES-UDS-47509 | 34.264602 | -5.232586 | 5.673 | 22.88 | 9.03 | 13.86 | 2062 | 7.54 |
RUBIES-EGS-27915 | 214.844229 | 52.789595 | 5.680 | 18.83 | 13.69 | 5.16 | 3060 | 7.69 |
RUBIES-EGS-37124 | 214.990977 | 52.916524 | 5.682 | 9.99 | 2.18 | 7.82 | 1610 | 7.20 |
CEERS-397 | 214.836197 | 52.882693 | 6.000 | 21.07 | 16.81 | 4.28 | 2483 | 7.49 |
RUBIES-EGS-49140 | 214.892248 | 52.877410 | 6.685 | 107.46 | 0.07 | 107.36 | 2782 | 8.30 |
RUBIES-UDS-807469 | 34.376139 | -5.310366 | 6.778 | 7.58 | 2.10 | 5.48 | 1978 | 7.40 |
Note. — The listed FWHM of the broad component is corrected for instrumental broadening. The stated errors on do not take into account the systematic uncertainties from the Reines et al. (2013) prescription. As noted in the text, RUBIES-EGS-50052 and CEERS-2782 are the same object observed independently by RUBIES and CEERS, and are thus marked with a (*). We attribute the differences in both line flux and FWHM to the extreme spatially dependent slit-loss seen in CEERS-2782.

### v2

@@ 3.2 Flux Calibration
While NIRSpec MSA flux calibration has improved significantly since the first data releases and pipelines (e.g. Schaerer et al., 2022; Taylor et al., 2022; Carnall et al., 2023; Trump et al., 2023), additional corrections to account for slit-losses/path-losses and overall photometric calibrations are still necessary. We use the NIRCam photometry described in §2.1 to account for these effects in our dataset. To do this, we first visibly inspect the 2D spectrum of each object for contamination. The spectra of some objects are contaminated by either positive flux from other sources in the same MSA row, or negative flux resulting from nodded background subtraction of sources in adjacent MSA rows. We visually inspect our sources and flag those that are impacted by either type of contamination. We find 11 sources with significant contamination, and 52 sources that are contamination free.
@@
For contamination-free sources, we apply a simple flux correction based on NIRCam photometry. For each source, we use the NIRCam filter transmission curves and G395M spectrum to measure synthetic photometry from the G395M spectrum for both the F356W and F444W NIRCam filters. We then select the filter that overlaps the detected broad-line and scale the flux and error spectra such that its broadband spectroscopic flux density matches the flux density of the broadband photometry. While using the higher S/N PRISM spectra (when available) would be preferable for this calibration, due to the different MSA configurations between the PRISM and G395M data, the measured fluxes between the two are not directly comparable. For the remaining 11 contaminated sources and three sources that fall outside the CEERS NIRCam footprint (see Figure 6), we instead apply the median corrective factor (1.28) from the 52 uncontaminated sources with NIRCam coverage to avoid any extreme calibration errors from the contamination.
After this correction, we re-fit the broad-lines using the emcee and above-mentioned broad-line model to produce final, flux-calibrated line fits and calibrated line measurements using the full calibrated line fit posterior distributions. We note that four of our objects (RUBIES-EGS-28812, RUBIES-EGS-42046, RUBIES-EGS-49140, and RUBIES-UDS-146995) visually show significant narrow absorption features on their H lines, similar to those exhibited in sources identified in Matthee et al. (2024) and Kocevski et al. (2024). When refitting these objects, we introduce a additional absorption component to our line model. We model this component as a Gaussian with a strictly negative amplitude and allow its line center and FWHM to vary freely. Notably, we observe blue-ward velocity offsets of the absorption features relative to the center of the H emission lines of km s-1, km s-1, km s-1, and km s-1, for RUBIES-UDS-146995, RUBIES-EGS-28812, RUBIES-EGS-42046, and RUBIES-EGS-49140, respectively. While the physical interpretation of these features is actively discussed in the literature (e.g., Matthee et al., 2024; Inayoshi & Maiolino, 2024), we defer a full analysis of these absorption features to future work.
@@
We apply a similar procedure to calibrate the PRISM spectra (when available) of the broad-line sample. Here, for a given source, we use the transmission curves of the F150W, F200W, F277W, F365W, F410M, and F444W filters to measure broadband fluxes for each filter. We then fit a first order polynomial to the ratio of the fluxes measured from the spectra and the fluxes reported in the broadband photometry as a function of the logarithm of the filter centers. We use this best fit line as a wavelength dependent multiplicative correction curve, and apply this correction to the spectral data. We repeat this process five times to ensure reasonable convergence between the spectrum and measured photometry. We use these calibrated PRISM spectra in §5 to measure optical continuum slopes.
@@ 3.3 Outflows
We next test our sample of broad-line objects for evidence of inflows/outflows to help distinguish true BLAGN from objects that have broad emission components due to strong kinematics in the ISM. For the 25 objects at redshifts , such that the [O iii]4959,5007 doublet falls within the G395M spectra, we examine the [O iii]4959,5007 doublet for evidence of broad lines.
@@
Here, we design two models for MCMC fitting. In the first model (the outflow model), we model [O iii]4959,5007 each as two Gaussians (one narrow and one broad). We enforce a common narrow and a common broad FWHM for both lines in the [O iii]4959,5007 doublet. In contrast to our H broad-line search, here we allow the line centers of the narrow and broad line components to vary to account for a velocity offset from an outflow. We fix the amplitude ratio of the [O iii] doublet lines to 1:2.98 for both the narrow and broad components. We also include a simple linear model for any faint continuum. In the second model (the BLAGN model), we exclude the broad components of the [O iii]4959,5007.
@@
We fit the spectra with both the outflow and BLAGN models each using 32 walkers and a 5000 step chain, of which we discard the first 3333 steps. Here, we also compute the Bayesian Information Criterion (BIC) for each model fit: , where is the number of free parameters in a fitted model, is the maximized likelihood function, and is the number of fitted data points. We find only one object (RUBIES-EGS-50052) with BICoutflow-BIC and a broad-line SNR, indicating significant broad [O iii]4959,5007 emission likely caused by a strong outflow. We show the single and double component fits to this object in Figure 3. This object exhibits an [O iii]4959,5007 broad component with a FWHM of km s-1, and an H broad component with a FWHM of km s-1. These FWHMs demonstrate the [O iii]4959,5007 doublet has a slightly lower velocity broad component than H, possibly due to kinematics independent of the BLAGN. Curiously, RUBIES-EGS-50052 is the source that was observed by both RUBIES and CEERS (as CEERS-2782). Its CEERS spectrum shows no trace of an outflow, but due to the aforementioned expected slit-losses for this source in CEERS, we expect that the absence of detected outflows may simply be due to the lower S/N of the object in the CEERS data relative to the RUBIES data. Alternatively, the outflow component may be spatially dependent, and may be better sampled by the RUBIES shutter than the CEERS shutter. While the exact nature of these spatially dependent effects is beyond the scope of this work, the difference in the FWHMs of the broad H emission and broad [O iii]emission from this source suggest a BLAGN signature, and we therefore keep RUBIES-EGS-50052 in our BLAGN sample. Nonetheless, if the broad H emission from RUBIES-EGS-50052 is indeed the result of an outflow and not an AGN, our large 62 object sample size ensures that the inclusion or exclusion of RUBIES-EGS-50052 will not significantly affect our results.
@@
Though we cannot perform this analysis for objects in our sample, the 24/25 objects that show clear BLAGN signatures that kinematically dominate any outflow velocity structure suggest that our full sample is likely significantly dominated by BLAGN, with low likelihood of containing a significant number of outflow dominated objects whose broad H emission is not driven by a central AGN. This result is consistent with the analysis performed in other collections of broad-line AGN detected by JWST at . Notably, Maiolino et al. (2024) also finds that outflows play a negligible role in broadening the lines of the Type-I AGN used in their analysis.
We also repeat our outflow test on a stack of all 25 objects. Here, we stack these object spectra using an inverse variance weighting. Our outflow test indicates that an [O iii]4959,5007 broad-component is very weakly preferred in the stacked spectrum with a BIC. However this broad component is only detected with a SNR=1.1, and is thus not significant. As such, we maintain our conclusion that our sample indeed contains BLAGN and is not significantly contaminated by objects with strong outflows.
@@
Beyond outflows, Baggen et al. (2024) propose that the broad-lines seen in high-redshift galaxies with JWST may instead be generated by highly-compact galaxies with intrinsically high virial velocities that can mimic the broad-lines typically associated with BLAGN. However, as 59/62 objects in our sample show both broad and narrow components in H, we expect that these are indeed BLAGN showing narrow emission from galaxy kinematics (or the AGN narrow line regions) and broad emission from the AGN, and are not the exotically dense galaxies proposed by Baggen et al. (2024). While the final object (RUBIES-EGS-49140) is analyzed in Wang et al. (2024b) and studied in Baggen et al. (2024) as a candidate for non-AGN broad-lines, we nonetheless keep this object in our BLAGN sample as the Baggen et al. (2024) model does not preclude this object from being a BLAGN, it simply offers a novel alternate explanation for its broad spectral features.
Source ID | R.A. | Decl. | FWHMHα,broad | |||||
[deg] | [deg] | [ erg s-1 cm-2] | [km s-1] | |||||
RUBIES-UDS-44043† | 34.241817 | -5.239436 | 3.499 | 31.53 | 8.33 | 23.25 | 2728 | 7.66 |
RUBIES-EGS-50812 | 214.845487 | 52.848281 | 3.519 | 12.07 | 5.85 | 6.24 | 1358 | 6.77 |
RUBIES-UDS-21944 | 34.469218 | -5.283563 | 3.526 | 13.69 | 2.14 | 11.54 | 1459 | 6.96 |
RUBIES-UDS-46566 | 34.293543 | -5.234588 | 3.538 | 124.88 | 74.50 | 50.39 | 747 | 6.67 |
RUBIES-UDS-154183 | 34.410749 | -5.129664 | 3.546 | 58.67 | 9.25 | 49.45 | 2234 | 7.64 |
RUBIES-UDS-153207 | 34.493112 | -5.130999 | 3.597 | 310.80 | 127.70 | 183.25 | 857 | 7.06 |
RUBIES-EGS-50522 | 214.855980 | 52.854661 | 3.614 | 4.61 | 1.34 | 3.28 | 1733 | 6.87 |
RUBIES-EGS-920 | 215.052344 | 52.884268 | 3.616 | 34.24 | 28.12 | 6.16 | 823 | 6.33 |
RUBIES-UDS-150323 | 34.417822 | -5.134842 | 3.618 | 7.51 | 2.91 | 4.61 | 2178 | 7.15 |
RUBIES-EGS-58237 | 214.850571 | 52.866030 | 3.651 | 63.78 | 14.34 | 49.47 | 2137 | 7.62 |
RUBIES-UDS-61627 | 34.238394 | -5.205775 | 3.654 | 8.40 | 4.36 | 4.04 | 1515 | 6.80 |
RUBIES-UDS-5496 | 34.405872 | -5.312951 | 3.655 | 13.30 | 6.42 | 6.94 | 1657 | 6.99 |
RUBIES-EGS-15825 | 215.079264 | 52.934252 | 3.666 | 72.08 | 7.29 | 64.80 | 3343 | 8.08 |
RUBIES-UDS-18302 | 34.233628 | -5.283850 | 3.698 | 21.30 | 6.32 | 15.38 | 1425 | 7.01 |
RUBIES-UDS-46885 | 34.291666 | -5.233788 | 3.730 | 25.73 | 15.46 | 10.31 | 709 | 6.32 |
RUBIES-UDS-146995‡ | 34.331043 | -5.139963 | 3.732 | 26.26 | 0.90 | 25.23 | 1403 | 7.11 |
RUBIES-EGS-34978 | 214.861690 | 52.818438 | 3.772 | 91.02 | 68.01 | 23.04 | 1485 | 7.15 |
RUBIES-EGS-19174 | 214.860840 | 52.784773 | 3.774 | 72.89 | 55.44 | 17.62 | 878 | 6.62 |
RUBIES-UDS-10036† | 34.381671 | -5.303742 | 3.806 | 9.46 | 3.12 | 6.34 | 1234 | 6.73 |
RUBIES-EGS-37032† | 214.849388 | 52.811824 | 3.850 | 18.00 | 5.09 | 13.08 | 1960 | 7.29 |
CEERS-11728 | 215.084870 | 52.970738 | 3.869 | 9.88 | 7.08 | 2.86 | 1191 | 6.53 |
RUBIES-UDS-22304 | 34.399170 | -5.283007 | 3.906 | 24.14 | 19.88 | 4.27 | 844 | 6.32 |
RUBIES-UDS-16053† | 34.367104 | -5.293524 | 3.952 | 8.18 | 2.13 | 6.06 | 1198 | 6.71 |
RUBIES-UDS-147411 | 34.360718 | -5.139081 | 3.966 | 8.42 | 3.33 | 5.11 | 1100 | 6.60 |
RUBIES-UDS-11721 | 34.411039 | -5.300780 | 3.978 | 33.51 | 23.75 | 9.78 | 1937 | 7.24 |
RUBIES-UDS-8895 | 34.363041 | -5.306108 | 3.982 | 120.33 | 67.12 | 53.25 | 1523 | 7.37 |
RUBIES-UDS-30969 | 34.296356 | -5.268672 | 4.000 | 51.14 | 40.29 | 10.92 | 1079 | 6.74 |
RUBIES-UDS-155916 | 34.317031 | -5.127611 | 4.098 | 7.93 | 0.16 | 7.72 | 1646 | 7.06 |
RUBIES-UDS-31747† | 34.223757 | -5.260245 | 4.130 | 16.16 | 4.85 | 11.32 | 1688 | 7.17 |
RUBIES-UDS-119957† | 34.268908 | -5.176722 | 4.149 | 12.83 | 4.23 | 8.61 | 1845 | 7.19 |
RUBIES-EGS-28812‡ | 214.924149 | 52.849050 | 4.222 | 17.59 | 7.24 | 10.37 | 2098 | 7.35 |
RUBIES-UDS-143683 | 34.316389 | -5.144678 | 4.226 | 15.94 | 11.95 | 4.05 | 1426 | 6.81 |
RUBIES-UDS-35974 | 34.331644 | -5.260593 | 4.367 | 41.29 | 33.18 | 8.16 | 1179 | 6.80 |
RUBIES-UDS-63139 | 34.230848 | -5.202607 | 4.435 | 5.45 | 2.53 | 2.92 | 1107 | 6.55 |
RUBIES-UDS-48507 | 34.284578 | -5.230702 | 4.468 | 19.68 | 11.22 | 8.47 | 762 | 6.43 |
CEERS-1244 | 215.240652 | 53.036041 | 4.478 | 50.54 | 16.55 | 34.02 | 2619 | 7.82 |
RUBIES-EGS-29489† | 215.022071 | 52.920786 | 4.543 | 13.48 | 2.21 | 11.32 | 2055 | 7.39 |
RUBIES-UDS-19484† | 34.232426 | -5.280654 | 4.656 | 3.76 | 1.73 | 2.06 | 1278 | 6.62 |
RUBIES-UDS-182791 | 34.213813 | -5.087050 | 4.718 | 40.12 | 8.20 | 31.91 | 3350 | 8.06 |
RUBIES-EGS-45599 | 214.831103 | 52.824124 | 4.862 | 7.57 | 4.68 | 2.90 | 1498 | 6.86 |
RUBIES-EGS-6411 | 215.109185 | 52.939770 | 4.880 | 7.26 | 4.96 | 2.32 | 1046 | 6.49 |
RUBIES-EGS-42232† | 214.886792 | 52.855381 | 4.954 | 23.59 | 4.48 | 19.14 | 2375 | 7.67 |
RUBIES-EGS-46985 | 214.805654 | 52.809497 | 4.963 | 16.06 | 9.46 | 6.67 | 1290 | 6.90 |
RUBIES-EGS-17146 | 214.949482 | 52.845415 | 5.000 | 36.80 | 21.82 | 15.01 | 763 | 6.60 |
RUBIES-EGS-50052 | 214.823454 | 52.830277 | 5.240 | 37.85 | 20.07 | 17.78 | 2129 | 7.58 |
CEERS-2782 | 214.823453 | 52.830281 | 5.240 | 26.41 | 14.47 | 11.95 | 1462 | 7.16 |
RUBIES-EGS-13872 | 215.132933 | 52.970705 | 5.262 | 7.67 | 4.45 | 3.31 | 1223 | 6.74 |
RUBIES-EGS-42046†‡ | 214.795368 | 52.788847 | 5.276 | 190.13 | 80.88 | 109.59 | 3146 | 8.30 |
RUBIES-UDS-970351† | 34.261900 | -5.105205 | 5.282 | 6.27 | 1.05 | 5.21 | 1657 | 7.11 |
RUBIES-EGS-926125† | 215.137081 | 52.988554 | 5.284 | 14.45 | 4.46 | 9.98 | 1888 | 7.36 |
RUBIES-EGS-60935† | 214.923373 | 52.925593 | 5.287 | 22.82 | 6.84 | 16.02 | 2144 | 7.57 |
RUBIES-UDS-29813† | 34.453355 | -5.270717 | 5.440 | 12.06 | 4.64 | 7.43 | 2047 | 7.39 |
RUBIES-UDS-172350† | 34.368951 | -5.103941 | 5.580 | 33.19 | 5.59 | 27.61 | 2068 | 7.68 |
CEERS-746† | 214.809142 | 52.868484 | 5.623 | 6.59 | 1.93 | 4.68 | 2021 | 7.29 |
CEERS-672† | 214.889677 | 52.832977 | 5.666 | 2.41 | 0.99 | 1.45 | 1433 | 6.75 |
RUBIES-UDS-19521 | 34.383672 | -5.287732 | 5.669 | 11.29 | 5.46 | 5.84 | 1682 | 7.18 |
RUBIES-UDS-47509† | 34.264602 | -5.232586 | 5.673 | 22.92 | 9.05 | 13.87 | 2063 | 7.54 |
RUBIES-EGS-37124† | 214.990977 | 52.916524 | 5.682 | 9.94 | 1.93 | 7.98 | 1566 | 7.18 |
RUBIES-UDS-139709† | 34.296002 | -5.149895 | 5.685 | 39.47 | 11.11 | 28.37 | 2552 | 7.88 |
CEERS-397 | 214.836197 | 52.882693 | 6.000 | 21.75 | 16.82 | 4.88 | 2727 | 7.60 |
RUBIES-UDS-174752 | 34.205808 | -5.100500 | 6.039 | 14.82 | 8.54 | 6.28 | 1741 | 7.26 |
RUBIES-EGS-49140†‡ | 214.892248 | 52.877410 | 6.685 | 116.81 | 4.94 | 111.71 | 2613 | 8.26 |
RUBIES-UDS-807469 | 34.376139 | -5.310366 | 6.778 | 7.72 | 2.18 | 5.53 | 2044 | 7.43 |
Note. — The listed FWHM of the broad component is corrected for instrumental broadening. The stated errors on do not take into account the systematic uncertainties from the Reines et al. (2013) prescription. We mark LRD sources with a (†), and we mark the four sources fit with an absorption feature with a (‡). As noted in the text, RUBIES-EGS-50052 and CEERS-2782 are the same object observed independently by RUBIES and CEERS, and are thus marked with a (*). We attribute the differences in both line flux and FWHM to the extreme spatially dependent slit-loss seen in CEERS-2782.
@@ 4 The BLAGN Sample
Here, we present our final sample of 62 H spectroscopically confirmed BLAGN. We show galleries of the fitted H broad-lines in Figure 4 and Figure 5, and detail the properties of these objects in Table 1. We show the positions of the BLAGN in the CEERS (EGS) and PRIMER-UDS fields, as well as the footprints of the CEERS and RUBIES NIRSpec MSA pointings in Figure 6. Three sources that lack NIRCam imaging coverage are visible in the upper left of the top panel and were targeted in the CEERS survey based on archival HST CANDELS imaging (Grogin et al., 2011; Koekemoer et al., 2011). The sample clearly exhibits a large diversity in H line shapes, [N II]6548,6583 line strengths, and redshifts. We analyze the properties of the sample below.